OpenStudy (anonymous):

(X+15) degrees that's the BD bisets and the other says (4x-45) degrees and it says ABC . Find the value of x

OpenStudy (anonymous):

,: Angle ABD= (x+15) degrees, and angle DBC= (4x-45) degrees. x+15 = 4x-45 -3x = -60 x = 20